OpenStudy (superhelp101):

1. If 13.4 grams of iron (III) oxide reacts with 12.3 grams of carbon monoxide to produce 8.23 g of pure iron, what are the theoretical yield and percent yield of this reaction? Be sure to show the work that you did to solve this problem. unbalanced equation: Fe2O3 + CO Fe + CO2

OpenStudy (superhelp101):

this is what i done so far: 1. The reaction balanced: Fe2O3 + 3 CO-->2 Fe + 3 CO2 2. Now, since the question says 13.4 grams of iron (III) oxide we have to Calculate the moles of Fe2O3 you have in 13.4 g. Which is 0.08391352648473714 3. Now, since the question above says 12.3 grams of carbon monoxide we have to calculate the moles of CO you have in 12.3 g. Which is 0.43912731479002215 @Abmon98

OpenStudy (abmon98):

If you have 2 different number of moles we use mole ratio to find out which is the limiting reagent. Mole ratio of Fe2O3 to CO 1:3 0.08:x x=0.08*3/1=0.24 moles of CO the calculated number of moles is less than the present number of moles so CO is in excess and Fe2O3 is the limiting reagent. Limiting reagent -The reactant in a chemical reaction that limits the amount of product that can be formed. Use mole ratio of CO to Fe 3:2 0.24:x moles of Fe 0.24*2/3=0.16 moles Number of moles=Mass(g)/ Molar Mass(g/mol) 8.928g=0.16*55.8=Mass Actual yield/Theoretical Yield (8.23/8.928)*100

OpenStudy (superhelp101):

thank you soo much this is super helpful!

OpenStudy (abmon98):

good to hear that and your welcome :)